consider the linear system
x+2y+3z=a
x+3y+8z=b
x+2y2z=c
where a,b and c are arbitrary constant.find all solutions of the system
1
Expert's answer
2020-04-28T14:36:53-0400

"\begin{cases}
 x+2y+3z=a \\
 x+3y+8z=b\\
x+2y+2z=c
\end{cases}"

"\begin{cases}
 x+2y+3z=a \\
 x+3y+8z=b\\
-z=c-a
\end{cases}"


"\begin{cases}
x+2y+3z=a\\
y+5z=b-a\\
z=a-c
\end{cases}"


"\begin{cases}
x+2y+3z=a\\
y+5(a-c)=b-a\\
z=a-c
\end{cases}"


"\begin{cases}
x+2y+3z=a\\
y=b-6a+5c\\
z=a-c

\end{cases}"


"\begin{cases}
x+2(b-6a+5c)+3(a-c)=a\\
y=b-6a+5c\\
z=a-c
\end{cases}"


"\begin{cases}
x=10a-2b-7c\\
y=b-6a+5c\\
z=a-c
\end{cases}"


Answer:

"x=10a-2b-7c"

"y=b-6a+5c"

"z=a-c"